# Low-Latency Event-Based Velocimetry for Quadrotor Control in a Narrow Pipe

### v1

@@ I Introduction
In our work, we go one step further and implement a sparse, low-latency, event-based smoke velocimetry method to measure the airflow in real time, then estimate the aerodynamic disturbances with a recurrent convolutional network and feed this information into an learning-based controller. This not only improves the hovering accuracy but also allows more controlled translational motion in the pipe cross-section. To the best of our knowledge, this is the first time closed-loop control of an aerial robot with real-time flow field measurements has been demonstrated. Event cameras are chosen for their unique ability to deliver low-latency, motion-blur-free data under difficult lighting conditions, without the complexity, cost, and bandwidth limitations of traditional high-speed imaging systems with pulsed lasersheet illumination.
@@
  • the first event-based smoke velocimetry method; it has a sub-millisecond processing latency and achieves a mean error of 0.35 m/s w.r.t. offline estimates obtained with the SoTA software PIVLab [17, 18, 19],
@@
  • the development of the first monocular event-based motion-capture system; it has a sub-millisecond processing latency and achieves millimeter accuracy while only using small IR-LEDs as active markers, and
@@ III-A Preliminary Considerations
This back-of-the-envelope calculation clearly shows three aspects: First, the flow estimation algorithm must run at very high update rates that are much faster than 100 Hz in order to be able to track features, since they can disappear as quickly as 10 ms. Second, obtaining per-pixel optical flow estimates is meaningless, as the observable smoke structures undergo appearance changes constantly. We need to evaluate the flow on larger scales. Third, the assumption of brightness constancy fundamental to many event-based methods is violated, as smoke structures can enter/exit the lightsheet. The scattered light can also illuminate the whole pipe, as it will contain some smoke.
@@ III-B SSD Template Matching with Quadratic Refinement
=$\mathcal{E}=\left\{e_{i}=\left(x_{i},y_{i},t_{i},p_{i}\right),~{}p_{i}\in\{-1,1\},~{}i=1\ldots N\,\right\},$ (1)
@@ III-B1 Template Matching
When dealing with low-contrast smoke structures, the events are sparse because they are triggered probabilistically by the structures. To avoid false minima in template matching, the image $I_{k}$ is blurred with a Gaussian kernel $G$ of size $7\times 7$ and a standard deviation of $\sigma_{\text{blur}}=1.75$. Also, to increase the number of events, we use overlapped event time-windows of length $n=3$:
@@ III-B2 Quadratic Refinement
The curvature of the quadratic surface is related to det $\text{det }\mathbf{A}$, and we discard all flow estimates with det $\text{det }\mathbf{A}\leq 0$ (maximum, not minimum) and $||\mathbf{u}_{\text{subpx}}||_{\infty}>1$ (minimum outside region). Additionally, we empirically find that det $\sqrt{\text{det }\mathbf{A}}$ is a good estimate of the confidence in the correctness of the flow estimate of a patch.
@@ III-B3 Resolution and Latency
Stacking the event frames, blurring the stacked image, and computing the cost function are implemented on the GPU using Nvidia’s NPPPlus SDK [54] and accelerated with graph tracing. For $P=11$ (121 patches) and all other parameters as given above, we obtain a runtime of $168\,\mathrm{\mu s}$ for stacking and blurring, and $344\,\mathrm{\mu s}$ for computing the cost function on an RTX 3090. Copying back the data to the CPU, finding the minimum, and refining it takes an additional $112\,\mathrm{\mu s}$, resulting in an overall runtime of $624\,\mathrm{\mu s}$, thus achieving sub-millisecond latency and making frequencies up to 1.6 kHz possible.
@@ III-C Wrench Estimation
Figure 2: Our network architecture used to estimate the aerodynamic disturbances. The ConvLSTM-based flow encoder outputs a latent representation of the flow configuration. This is concatenated with the position of the quadrotor and fed into an MLP to estimate the aerodynamic disturbances.
@@
The flow input is processed using a convolutional Long Short-Term Memory (ConvLSTM) network. This module enables the network to capture both spatial and temporal patterns within the flow field and smooth out noisy flow measurement vectors. The output of the ConvLSTM is a compact latent representation. This latent vector is then concatenated with the position input and passed to a two-layer perceptron (MLP), which outputs estimates of the horizontal disturbance force $f_{y}^{d}$, vertical disturbance force $f_{z}^{d}$, and roll torque disturbance $\tau_{x}^{d}$ acting on the vehicle.
A significant challenge in estimating aerodynamic disturbances is the presence of experimental biases. Such biases arise from physical asymmetries in the quadrotor (e.g., unbalanced mass distribution, e.g., off-center battery placement). While these biases are approximately constant within a single experiment, they can vary across different trials and corrupt the learned disturbance signal.
To address this, we introduce a parallel bias estimation module, implemented as a lightweight MLP, which is trained jointly with the primary disturbance network. This is only used during training and discarded afterwards. The bias estimation module only has access to a single input, the index of the experiment (simply an increasing sequence number), and thus can only estimate an additive, per-experiment bias. The total loss function $\mathcal{L}$ used for training combines two terms: a mean squared error (MSE) loss between the bias-corrected, predicted disturbance $\hat{d}+\hat{b}$, and the ground-truth disturbance $d_{\text{gt}}$; and an L2 regularization term on the estimated bias.
@@
The regularization ensures that the bias network only compensates when necessary, allowing the model to disentangle systematic biases from dynamic flow-induced effects.
@@ IV Event-Based Monocular Pose Estimation
Having addressed the topic of aerodynamic disturbance estimation in the previous section, we now focus on the next challenge: state estimation. A monocular event-based motion-capture system with active markers is ideally suited for this task. Figure 3 shows the developed millisecond-latency, millimeter-accuracy motion capture system, which takes inspiration from [48]. We first give an overview of the whole system and then describe the two key aspects enabling robust, accurate, and low-latency pose estimation: the novel SDTV (Signed Delta-Time Volume) event representation enabling a processing latency of $273\,\mathrm{\mu s}$ and our particle-swarm optimization, which optimally sets the contrast thresholds, biases, and filter parameters of the event camera. Details on the electrical design of the active LED markers and the pose estimation accuracy can be found in Appendix A and B.
@@ IV-B2 Latency
On a laptop running at a 2.9 GHz CPU clock and a VGA-resolution event camera, the processing latency incurred by finding the blinking LEDs, identifying their frequency, and clustering them is only $144\,\mathrm{\mu s}$. The particle filter used to track the centroids of the blinking LEDs and the SqPnP pose estimation add another $129\,\mathrm{\mu s}$ latency, resulting in a total processing latency of $t_{\text{proc}}=273\,\mathrm{\mu s}$. Consequently, we are able to truly take advantage of the low latency of the event camera.
Note that due to the slowest LED blinking only at ${f_{\text{min}}=1700\,\mathrm{Hz}}$, increasing the processing frequency beyond $f_{\text{min}}/2$ does not make sense, as at least one full LED cycle must be captured to identify the frequency. Thus, the overall worst-case latency is $2/f_{\text{min}}+t_{\text{proc}}=1.45\,\mathrm{ms}$.
@@ IV-C Automatic Event Camera Tuning
Consider a scenario where the event-camera and an object with $K$ markers with known frequencies ${f_{k},~{}k=1\ldots K}$ are statically placed. The pixel coordinates of each LED center are given by $(x_{k}^{c},y_{k}^{c})$. At a pixel $(x,y)$ illuminated by marker $k$, an event-stream of duration $t$ with events $n=1\ldots N$ ideally contains $t\,f_{k}$ positive $p=1$ and negative $p=-1$ events. We define the per-pixel event ratio $\alpha_{k}(x,y)^{+}$ for positive and $\alpha_{k}(x,y)^{-}$ for negative events as
@@
TABLE I: Default and Optimal Parameters for Gen3
@@
We employ PSO [20] with 100 particles to jointly optimize the biases $\theta$ of the event camera to minimize the above cost function $J$. We set $t=0.1\,\mathrm{s}$ and after 40 to 60 iterations (around 10 min time), the algorithm converges with an overall cost of $J^{*}=0$, i.e., no events are missed and not more than 50 % of the transitions trigger two events. The optimal parameters for a Prophesee Gen3 are reported in Tab. I.
@@ V Disturbance-Aware Quadrotor Controller
With both disturbance estimation and state estimation addressed, the last component of our system is the learning-based controller which takes the real-time disturbance measurements into account. Following the demonstrated success of reinforcement learning (RL) in enabling robust quadrotor flight in challenging real-world conditions [1, 61], we choose an RL-based approach in this work.
@@ V-A LSTM-PPO Policy Architecture
where ${\mathbf{\hat{p}}\in\mathbb{R}^{3}}$ is the estimated quadrotor position, ${\mathbf{\hat{R}}_{12}(:)\in\mathbb{R}^{6}}$ denotes the vectorized first two columns of the rotation matrix representing estimated orientation, and ${\mathbf{p}_{\text{target}}\in\mathbb{R}^{3}}$ is the goal position. The term ${\mathbf{a}_{t-1}\in\mathbb{R}^{3}}$ denotes the action applied in the previous timestep. The estimated disturbance wrench is given by ${[\hat{f}_{y}^{d},\hat{f}_{z}^{d},\hat{\tau}_{x}^{d}]\in\mathbb{R}^{3}}$, and $\zeta\in\{0,1\}$ is a binary indicator flag that signals whether valid disturbance measurements are available at the current timestep (e.g., is smoke currently injected into the pipe). This mechanism enables the policy to learn how to handle both nominal conditions (without flow sensing) and flow-aware control when the estimator is active.
@@ V-B Training in Simulation
The unsteady, turbulent component is modeled by analyzing real-world data. First, we subtract the mean, position-dependent disturbance from measured wrench data to isolate the residual. The power spectral density (PSD) of this residual is then estimated and used to fit an autoregressive (IIR) filter via the Yule-Walker method. During simulation, temporally correlated noise is generated by filtering Gaussian white noise through this IIR model, reproducing the correct spectral characteristics of the real-world disturbances.
@@ V-C Training Details
An episode is considered successful if the quadrotor reaches the target within a threshold ($||\mathbf{p}-\mathbf{p}_{\text{target}}||<0.01\,\mathrm{m}$) and stabilizes with low translational and rotational velocities ($||\mathbf{v}||<0.1\,\mathrm{m/s}$, $||\boldsymbol{\omega}||<0.1\,\mathrm{rad/s}$).
@@ VI Experimental Setup
TABLE II: Specifcations of the Quadrotor
@@
To perform the experiments, a small and lightweight quadrotor is equipped with the active LED markers detailed in Sec. A. The specifications of the vehicle are summarized in Tab. II. The quadrotor’s tip-to-tip length of 19.5 cm corresponds to half the pipe diameter. With a thrust-to-weight ratio of about 7, the quadrotor has enough thrust to avoid saturation of any motor even in challenging scenarios.
@@ VII-A1 Symmetric Configuration
First, we consider the case where the quadrotor hovers in the center (e.g. ${y=0}$ and ${z=R/2}$), as shown in Fig. 6. Since the setup is symmetric w.r.t. to $y=0$, the flow can develop such that the downwash from the left propellers gets recirculated on the left in a clockwise vortex, and the downwash from the right propellers gets recirculated to the right in a counterclockwise vortex. This flow configuration is shown in Fig. 6a.

### v2

@@ I Introduction
In our work, we go one step further and implement a sparse, low-latency, event-based smoke velocimetry method to measure the airflow in real time, then estimate the aerodynamic disturbances with a recurrent convolutional network and feed this information into an learning-based controller. This not only improves the hovering accuracy but also allows more controlled translational motion in the pipe cross-section. To the best of our knowledge, this is the first time closed-loop control of an aerial robot with real-time flow field measurements has been demonstrated. Event cameras are chosen for their unique ability to deliver low-latency, motion-blur-free data under difficult lighting conditions, without the complexity, cost, and bandwidth limitations of traditional high-speed imaging systems with pulsed lasersheet illumination. Following standard practice in experimental fluid dynamics, velocimetry is performed using an external camera rather than an onboard sensor, which simplifies the measurement process and avoids measurement-induced flow perturbations.
@@
  • the first event-based smoke velocimetry method; it has a sub-millisecond processing latency, runs on embedded computers, and achieves a mean error of 0.35 m/s w.r.t. offline estimates obtained with the SoTA software PIVLab [17, 18, 19],
@@
  • the development of the first monocular event-based motion-capture system; it has a sub-millisecond processing latency, runs on embedded computers, and achieves millimeter accuracy while only using small IR-LEDs as active markers, and
@@ III-A Preliminary Considerations
Figure 2: Event frames of the turbulent airflow inside the pipe. The frames are spaced 10 ms apart and polarity-integrated over a 6 ms window. Characteristic smoke structures change appearance on very short timescales, necessitating a high velocimetry update rate and the use of larger spatial patches rather than dense, pixel-level flow estimates.
This back-of-the-envelope calculation clearly shows three aspects: First, the flow estimation algorithm must run at very high update rates that are much faster than 100 Hz in order to be able to track features, since they can disappear as quickly as 10 ms. Second, obtaining per-pixel optical flow estimates is meaningless, as the observable smoke structures undergo appearance changes constantly. We need to evaluate the flow on larger scales. Third, the assumption of brightness constancy fundamental to many event-based methods is violated, as smoke structures can enter/exit the lightsheet. The scattered light can also illuminate the whole pipe, as it will contain some smoke. Figure 2 exemplarily shows four event frames spaced 10 ms apart to illustrate the points above.
@@ III-B SSD Template Matching with Quadratic Refinement
$\mathcal{E}=\left\{e_{i}=\left(x_{i},y_{i},t_{i},p_{i}\right),~p_{i}\in\{-1,1\},~i=1\ldots N\,\right\},$ (1)
@@ III-B1 Template Matching
When dealing with low-contrast smoke structures, the events are sparse because they are triggered probabilistically by the structures. To increase the number of events, we use overlapped event time-windows of length $n=3$, resulting in event-frames as shown in Fig. 2. To avoid false minima in template matching, the image $I_{k}$ is blurred with a Gaussian kernel $G$ of size $7\times 7$ and a standard deviation of $\sigma_{\text{blur}}=1.75$:
@@ III-B2 Quadratic Refinement
The curvature of the quadratic surface is related to $\text{det }\mathbf{A}$, and we discard all flow estimates with $\text{det }\mathbf{A}\leq 0$ (maximum, not minimum) and $\mathbf{u}_{\text{subpx}}||_{\infty}>1$ (minimum outside region). Additionally, we empirically find that $\sqrt{\text{det }\mathbf{A}}$ is a good estimate of the confidence in the correctness of the flow estimate of a patch.
@@ III-B3 Resolution and Latency
TABLE I: Runtimes of the EBSV Algorithm
Embedded Runtime [ms] Workstation (CPU + GPU) Runtime [ms]
Jetson Orin NX 4.98 i9 8950HK + P3200 1.58
Jetson Thor 1.09 i9 12900K + RTX 3090 0.44
9950X3D + RTX 5090 0.33
Stacking the event frames, blurring the stacked image, and computing the cost function are implemented on the GPU using Nvidia’s NPPPlus SDK [54] and accelerated with graph tracing. We evaluate the performance of our EBSV-algorithm on both embedded systems and general-purpose computers. The runtimes for $P=11$ (121 patches) and all other parameters as given above are summarized in Tab. I. The credit-card sized Orin NX is already capable of achieving 200 Hz and more powerful embedded systems such as the Nvidia Jetson Thor reach close to 1 kHz. Workstations can run the flow estimator in real-time at over 500 Hz, with a modern GPU yielding sub-millisecond latency.
@@ III-C Wrench Estimation
Figure 3: Our network architecture used to estimate the aerodynamic disturbances. The ConvLSTM-based flow encoder outputs a latent representation of the flow configuration. This is concatenated with the position of the quadrotor and fed into an MLP to estimate the aerodynamic disturbances. During training we also train with an additive bias which is only a function of the experiment index and accounts for experimental biases, e.g. battery placement.
@@
The flow input is processed using a convolutional Long Short-Term Memory (ConvLSTM) network. This module enables the network to capture both spatial and temporal patterns within the flow field and smooth out noisy flow measurement vectors. We opted for the ConvLSTM architecture over more recent video-transformers [55] because their spatio-temporal attention and frame batching makes them less suited for sequential, real-time applications. The output of the ConvLSTM is a compact latent representation. This latent vector is then concatenated with the position input and passed to a two-layer perceptron (MLP), which outputs estimates of the horizontal disturbance force $f_{y}^{d}$, vertical disturbance force $f_{z}^{d}$, and roll torque disturbance $\tau_{x}^{d}$ acting on the vehicle.
A significant challenge in estimating aerodynamic disturbances is the presence of experimental biases. Such biases arise from physical asymmetries in the quadrotor (e.g., unbalanced mass distribution due to off-center battery placement). While these biases are approximately constant within a single experiment, they can vary across different trials and corrupt the learned disturbance signal.
To address this, we introduce a parallel bias estimation module, implemented as a lightweight MLP, which is trained jointly with the primary disturbance network as shown in Fig. 3. This is only used during training and discarded afterwards. The bias estimation module only has access to a single input, the index of the experiment (simply an increasing sequence number), and thus can only estimate an additive, per-experiment bias. The total loss function $\mathcal{L}$ used for training combines two terms: a mean squared error (MSE) loss between the bias-corrected, predicted disturbance $\hat{d}+\hat{b}$, and the ground-truth disturbance $d_{\text{gt}}$; and an L2 regularization term on the estimated bias.
@@
The regularization ensures that the bias network only compensates when necessary, allowing the model to disentangle systematic biases from dynamic flow-induced effects. Compared to bias-mitigation strategies that operate by enforcing symmetry in the estimated wrenches, our approach has the advantage to be applicable to arbitrarily shaped pipes and other drones as it makes no assumptions on the cause of the bias.
@@ IV Event-Based Monocular Pose Estimation
Having addressed the topic of aerodynamic disturbance estimation in the previous section, we now focus on the next challenge: state estimation. A monocular event-based motion-capture system with active markers is ideally suited for this task. Figure 4 shows the developed millisecond-latency, millimeter-accuracy motion capture system, which takes inspiration from [48]. We first give an overview of the whole system and then describe the two key aspects enabling robust, accurate, and low-latency pose estimation: the novel SDTV (Signed Delta-Time Volume) event representation enabling a processing latency of $273\,\mathrm{\mu s}$ on a laptop and our particle-swarm optimization, which optimally sets the contrast thresholds, biases, and filter parameters of the event camera. Details on the electrical design of the active LED markers and the pose estimation accuracy can be found in App. A and B.
@@ IV-B2 Latency
To fully take advantage of the low latency of the event camera, the processing must fast. In Tab. II we show the performance of our approach on different devices, ranging from embedded devices to desktop workstations. Notably, even on an old Raspberry Pi 4B we are able to run our motion-capture at 300 Hz and on modern Nvidia boards, speeds above 1 kHz are easily achieved. Consequently, we are able to truly leverage the low latency of the event camera and can also deploy the system on embedded hardware.
Looking at the breakdown of the overall runtime we note that the relative runtime of the SDTV-based LED detection and frequency estimation is much larger on the older ARM-cortex based systems (Raspberry Pi 4B and Jetson Orin NX). A possible explanation for this is the processor cache size which is 1 MB on the Raspberry Pi 4B and 4 MB on the Orin NX. At VGA resolution and $D=16$ the SDTV volume is 9.4 MB in size and thus does not fit into the cache. Frequent cache misses negatively affect the runtime, a problem that the Thor and other CPUs do not have, as their L3-caches are large enough to fit the entire SDTV volume.
TABLE II: Runtimes of the Event-based Pose-Estimation
Device Runtimes [ms] (% of total runtime)
Total Detect LED Filter SqPnP
Embedded Raspberry Pi 4B 3.22 2.48 (77%) 0.40 (12%) 0.34 (11%)
Jetson Orin NX 1.29 0.92 (71%) 0.23 (18%) 0.14 (11%)
Jetson Thor 0.61 0.38 (62%) 0.14 (23%) 0.09 (14%)
Computer Intel i9 8950HK 0.27 0.14 (52%) 0.07 (26%) 0.06 (22%)
Intel i9 12900K 0.17 0.09 (53%) 0.05 (29%) 0.03 (18%)
AMD 9950X3D 0.14 0.06 (43%) 0.04 (29%) 0.04 (28%)
Note that due to the slowest LED blinking only at $f_{\text{min}}=1700\,\mathrm{Hz}$}, increasing the processing frequency beyond $f_{\text{min}}/2$ does not make sense, as at least one full LED cycle must be captured to identify the frequency. Thus, the overall worst-case latency in this work is considered $2/f_{\text{min}}+t_{\text{proc}}$, e.g. $1.45\,\mathrm{ms}$ for the laptop workstation.
@@ IV-C Automatic Event Camera Tuning
Consider a scenario where the event-camera and an object with $K$ markers with known frequencies $f_{k},~k=1\ldots K$} are statically placed. The pixel coordinates of each LED center are given by $(x_{k}^{c},y_{k}^{c})$. At a pixel $(x,y)$ illuminated by marker $k$, an event-stream of duration $t$ with events $n=1\ldots N$ ideally contains $t\,f_{k}$ positive $p=1$ and negative p=1p=-1 events. We define the per-pixel event ratio $\alpha_{k}(x,y)^{+}$ for positive and $\alpha_{k}(x,y)^{-}$ for negative events as
@@
TABLE III: Default and Optimal Parameters for Gen3
@@
We employ PSO [20] with 100 particles to jointly optimize the biases $\theta$ of the event camera to minimize the above cost function $J$. We set $t=0.1\,\mathrm{s}$ and after 40 to 60 iterations (around 10 min time), the algorithm converges with an overall cost of $J^{*}=0$, i.e., no events are missed and not more than 50 % of the transitions trigger two events. The optimal parameters for a Prophesee Gen3 are reported in Tab. III. The very high contrast thresholds also effectively eliminate spurious events triggered by the moving smoke and strongly attenuate noise. Measurements show a noise event-rate of only 0.7 ev/ms, compared to 587 ev/ms with the blinking LEDs. The SDTV representation and subsequent processing fully prevent any spurious marker detections due to noise.
@@ V Disturbance-Aware Quadrotor Controller
With both disturbance estimation and state estimation addressed, the last component of our system is the learning-based controller which takes the real-time disturbance measurements into account. Following the demonstrated success of reinforcement learning (RL) in enabling robust quadrotor flight in challenging real-world conditions [1, 62], we choose an RL-based approach in this work. The RL-controller outputs collective thrust and body-rate (CTBR) commands [63], which are tracked by a low-level PID controller onboard the drone. During deployment, the integral term of the low-level controller compensates for static torque biases, effectively masking them from the RL-controller.
@@ V-A LSTM-PPO Policy Architecture
where $\mathbf{\hat{p}}\in\mathbb{R}^{3}$} is the estimated quadrotor position, $\mathbf{\hat{R}}_{12}(:)\in\mathbb{R}^{6}$} denotes the vectorized first two columns of the rotation matrix representing estimated orientation, and $\mathbf{p}_{\text{target}}\in\mathbb{R}^{3}$} is the goal position. The term $\mathbf{a}_{t-1}\in\mathbb{R}^{3}$} denotes the action applied in the previous timestep. The estimated disturbance wrench is given by $\hat{f}_{y}^{d},\hat{f}_{z}^{d},\hat{\tau}_{x}^{d}]\in\mathbb{R}^{3}$}, and $\zeta\in\{0,1\}$ is a binary indicator flag that signals whether valid disturbance measurements are available at the current timestep (e.g., is smoke currently injected into the pipe). This mechanism enables the policy to learn how to handle both nominal conditions (without flow sensing) and flow-aware control when the estimator is active.
@@ V-B Training in Simulation
The unsteady, turbulent component is modeled by analyzing real-world data. First, we subtract the mean, position-dependent disturbance from measured wrench data to isolate the residual. The power spectral density (PSD) of this residual is then estimated and used to fit an autoregressive (IIR) filter via the Yule-Walker method. During simulation, temporally correlated noise is generated by filtering Gaussian white noise through this IIR model, reproducing the correct spectral characteristics of the real-world disturbances. To validate our simulation approach in the time-domain, we treat the real-world and simulated disturbances as samples from probability distributions and compute the Wasserstein distance $W_{1}$ between them. For the force disturbances along the $y$ and $z$ axes, we obtain distances of 7.3 mN and 11.6 mN, respectively. For the roll-torque disturbances, the distance is 0.56 mNm, indicating close agreement between simulation and experiment.
@@ V-C Training Details
An episode is considered successful if the quadrotor reaches the target within a threshold ($\mathbf{p}-\mathbf{p}_{\text{target}}||<0.01\,\mathrm{m}$) and stabilizes with low translational and rotational velocities ($\mathbf{v}||<0.1\,\mathrm{m/s}$, $\boldsymbol{\omega}||<0.1\,\mathrm{rad/s}$).
@@ VI Experimental Setup
TABLE IV: Specifcations of the Quadrotor
@@
$r_{p}$$l$
To perform the experiments, a small and lightweight quadrotor is equipped with the active LED markers detailed in Sec. A. The specifications of the vehicle are summarized in Tab. IV. The quadrotor’s tip-to-tip length of 19.5 cm corresponds to half the pipe diameter. With a thrust-to-weight ratio of about 7, the quadrotor has enough thrust to avoid saturation of any motor even in challenging scenarios.
@@ VII-A1 Symmetric Configuration
First, we consider the case where the quadrotor hovers in the center (e.g. y=0{y=0} and z=R/2{z=R/2}), as shown in Fig. 7. Since the setup is symmetric w.r.t. to $y=0$, the flow can develop such that the downwash from the left propellers gets recirculated on the left in a clockwise vortex, and the downwash from the right propellers gets recirculated to the right in a counterclockwise vortex. This flow configuration is shown in Fig. 7a.